OpenStudy (automaticloveletter):

Line R is represented by the following equation: x + y = 2 Which equation completes the system that is satisfied by the solution (1, 1)? 2x + y = 2 4x − 2y = 2 2x − 2y = 2 x + y = 4

OpenStudy (sdfgsdfgs):

(1,1) is the solution so it must satisfy the two equations: 1st one x + y =2 and 1 + 1 =2 so that is good. With me so far?

OpenStudy (automaticloveletter):

yes

OpenStudy (sdfgsdfgs):

now put in x=1 and y=1 into the choices to see which eqn is good :)

OpenStudy (automaticloveletter):

b..?

OpenStudy (sdfgsdfgs):

u didnt put letters in front of ur choices, @automaticloveletter :P but yes, b is the right ans!

OpenStudy (automaticloveletter):

i did it in my head lol but thanks!

OpenStudy (sdfgsdfgs):

welcome :)

OpenStudy (automaticloveletter):

@sdfgsdfgs can you help me with some more?

OpenStudy (sdfgsdfgs):

sure @automaticloveletter just post the prob :)

OpenStudy (automaticloveletter):

Based on the graph, which statement is correct about the solution to the system of equations for lines A and B? (4, 4) is the solution to line A but not to line B. (4, 4) is the solution to both lines A and B. (5, 1) is the solution to line B but not to line A. (0, 2) is the solution to both lines A and B.

OpenStudy (sdfgsdfgs):

is there a graph?

OpenStudy (automaticloveletter):

http://prntscr.com/7kym0l

OpenStudy (sdfgsdfgs):

OK, read the choices n tell me what u think plz :P

OpenStudy (automaticloveletter):

b?

OpenStudy (sdfgsdfgs):

correct :)

OpenStudy (automaticloveletter):

Which description best describes the solution to the following system of equations? y = −x + 4 y = 3x + 3 Line y = −x + 4 intersects the line y = 3x + 3. Lines y = −x + 4 and y = 3x + 3 intersect the x-axis. Lines y = −x + 4 and y = 3x + 3 intersect the y-axis. Line y = −x + 4 intersects the origin.

OpenStudy (sdfgsdfgs):

@automaticloveletter u should know the ans already :P

OpenStudy (automaticloveletter):

i'm very clueless /: a?

OpenStudy (sdfgsdfgs):

@automaticloveletter u said clueless n yet got the right ans :P

OpenStudy (automaticloveletter):

cuz i am xD

OpenStudy (automaticloveletter):

A pair of equations is shown below. x + y = 5 y = one halfx + 2 If the two equations are graphed, at what point do the lines representing the two equations intersect? (2, 5) (5, 2) (2, 3) (3, 2)

OpenStudy (sdfgsdfgs):

this one u actually have to solve the solution....feel like do it @automaticloveletter ?